# Kerr enhanced optomechanical entanglement via reservoir design

### v1

@@ Abstract
Quantum entanglement is a crucial resource in quantum technologies, enabling advancements in quantum computing, quantum communication, and quantum precision measurement. Here, we propose a method to enhance optomechanical entanglement by introducing an optical Kerr nonlinear medium and a squeezed vacuum reservoir of the optomechanical cavity. By performing the displacement and squeezing transformations, the system can be reduced to a standard linearized optomechanical system with normalized driving detuning and linearized-coupling strength, in which the optical and mechanical modes are, respectively, coupled to an optical vacuum bath and a mechanical heat bath. We focus on the entanglement generation in the single stable regime of the system. By evaluating the steady-state logarithm negativity, we find that the optomechanical entanglement can be enhanced within a wide range of the Kerr constant. In addition, the Kerr nonlinearity can extend the stable region, enabling considerable entanglement generation in the blue sideband-parameter region. We also investigate the dependence of the entanglement generation on the average thermal phonon occupation of the mechanical bath and the optical driving amplitude. It is found that the presence of the Kerr nonlinearity allows the generation of optomechanical entanglement even when the thermal phonon occupation of the mechanical bath is as high as 3000. Our findings will provide valuable insights into enhancing fragile quantum resources in quantum systems.
@@ I Introduction
Cavity optomechanics primarily investigates the radiation-pressure interaction between optical modes and mechanical modes, along with the novel quantum effects and applications arising from this interaction [1, 2, 3]. The study of cavity optomechanics holds dual significance in both fundamental theoretical research [4, 5] and cutting-edge applied research [6]. Not only does it advance the fundamental theoretical exploration of quantum theory, including macroscopic quantum behaviors [7, 8] and quantum-classical boundary [4, 9], but it also advances the applied research value in frontier quantum technologies, especially concerning quantum precision measurement [10, 11, 12] and quantum sensing [13, 14, 15]. In cavity optomechanics, the optomechanical entanglement holds significant research meaning for both aspects [16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27], because quantum entanglement is located at the kernel position of quantum theory [28, 29], and it is an important resource in quantum information science [30].
In general, it is expected to generate strong entanglement for implementing various quantum information processing tasks. In cavity optomechanics, various strategies have been proposed, such as reservoir engineering techniques [31, 32, 33], multimode coupling and collective effects [34, 35, 36], the injection of squeezed light [37], and the quantum interference effect [38, 39]. Usually, quantum entanglement induced by interaction can be enhanced via increasing the coupling strength. For linearized optomechanical interaction, the coupling strength can be enhanced by increasing the driving strength. However, the large driving amplitude is subject to the unstability issue [40]. This motivates us to explore other methods to enhancing the optomechanical interaction.
@@ II Physical model
Figure 1: Schematic of the Kerr cavity optomechanical system consisting of a mechanical mode $b$ optomechanically coupled to a optical mode $a$ containing a Kerr nonlinear medium with nonlinear constant $\chi$. The $g_{0}$ is the single-photon optomechancial coupling strength. The cavity field and the mechanical resonator are, respectively, contacted to a squeezed vacuum reservoir (with decay rate $\kappa_{a}$, squeezing parameter $r_{e}$, and reference phase $\theta_{e}$) and a heat bath (with decay rate $\kappa_{b}$ and average thermal phonon occupation $\bar{n}_{th}$). In addition, the cavity field is driven by a monochromatic field with driving amplitude $\Omega$ and frequency $\omega_{d}$.
The system under consideration is a Kerr cavity optomechanical system, which is formed by an optomechanical cavity containing a Kerr nonlinear medium. We consider the adiabatic regime of the cavity field, and hence focus on a single field mode in the cavity. The cavity field is coupled to the mechanical mode through the radiation-pressure interaction. The cavity is driven by a monochromatic field, and it is connected to a squeezed vacuum bath, while the mechanical mode is coupled to a heat bath. In a rotating frame defined by the unitary transformation operator $\mathrm{exp}(-i\omega_{d}\hat{a}^{{\dagger}}\hat{a}t)$, with $\omega_{d}$ being the driving frequency, the Hamiltonian of the Kerr cavity optomechanical system reads ($\hbar=1$)
$~{}\hat{H}_{I}$ $=$ $\Delta_{c}\hat{a}^{\dagger}\hat{a}+\omega_{m}\hat{b}^{\dagger}\hat{b}-g_{0}\hat{a}^{\dagger}\hat{a}(\hat{b}^{\dagger}+\hat{b})$ (1)
@@
To include the dissipation in the system, we assume that the cavity mode is connected to a squeezed vacuum reservoir with the central frequency $\omega_{d}$, and the mechanical mode is coupled to a heat bath. Here, the squeezed vacuum reservoir is introduced to counteract the thermal noise caused by the two-photon terms, which is induced by the Kerr nonlinearity under the linearization. In the Markovian-dissipation regime, the evolution of the system is governed by the quantum master equation
@@ III Enhanced optomechanical interaction
where $\hat{D}_{a}(\alpha)=\exp(\alpha\hat{a}^{\dagger}-\alpha^{*}{\hat{a}})$ and ${\hat{D}}_{b}(\beta)=\exp(\beta{\hat{b}}^{\dagger}-\beta^{*}{\hat{b}})$ are, respectively, the displacement operators of modes $a$ and $b$, with the time-dependent displacement amplitudes $\alpha(t)$ and $\beta(t)$. In the displacement representation, the quantum master equation (2) becomes
@@
In the derivation of Eq. (III), we have eliminated the linear term of the operators $\hat{a}$ ($\hat{a}^{\dagger}$) and $\hat{b}$ ($\hat{b}^{\dagger}$) by setting their coefficients to be zero. In this way, we obtain the equations of motion of the displacement amplitudes $\alpha(t)$ and $\beta(t)$ as follows,
@@
Here, $\hat{\rho}_{sd}$ is the density matrix of the system in the squeezed representation, and $\hat{S}_{a}(\xi)=\exp[(\xi^{\ast}\hat{a}^{2}-\xi\hat{a}^{\dagger 2})/2]$ is the squeezing operator with $\xi=r\exp(i\varphi)$, where $r$ is the squeezing parameter and $\varphi$ is the reference phase for the squeezed field. In the squeezed representation, the quantum master equation is given by
@@
where $\alpha_{r}$ and $\alpha_{i}$ are, respectively, the real and imagine parts of $\alpha_{ss}$, i.e., $\alpha_{ss}=\alpha_{r}+i\alpha_{i}$.
@@
×$\!\times(\sinh r_{e}\sinh r\!+e^{i(\varphi\!-\theta_{e})}\cosh r_{e}\cosh r).$ (14b)
@@
Below, we analyze how to obtain the solution of $r_{e}$ and $\theta_{e}$. In principle, under the given $r$ and $\varphi$, we can analyze the parameters $N_{ss}$ and $M_{ss}$ as functions of $\theta_{e}$ and $r_{e}$. By analyzing Eqs. (14), we find that $N_{ss}=0$ and $M_{ss}=0$ when $\theta_{e}=\varphi+n\pi$ $(n=1,3,5,\dots)$ and $r_{e}=r$. It should be emphasized that $N_{ss}=0$ and $M_{ss}=0$ imply that both the thermal noise and the squeezing vacuum noise have been suppressed completely. This working point is very important to observe the optomechanical effects at the single-photon level [41]. This feature is also the motivation for introducing the squeezing vacuum reservoir, i.e., using a well-designed squeezing vacuum bath to suppress the thermal noise and excitations caused by the two-photon driving.
Figure 2: The normalized detuning $\Delta_{sd}/\omega_{m}$ versus the scaled bare driving detuning $\Delta_{c}/\omega_{m}$ and the scaled Kerr constant $\chi/\omega_{m}$ when the cavity-field decay rate $\kappa_{a}$ takes different values: (a) $\kappa_{a}/\omega_{m}=0.5$, (b) $\kappa_{a}/\omega_{m}=0.8$, (c) $\kappa_{a}/\omega_{m}=1.2$, and (d) $\kappa_{a}/\omega_{m}=1.5$. Other parameters used are $g_{0}/\omega_{m}=0.005$, $\kappa_{b}/\omega_{m}=10^{-5}$, $\Omega/\omega_{m}=50$, and $\bar{n}_{th}=0$. Here, we also show the surfaces corresponding to $\Delta_{sd}/\omega_{m}=1$ and $\Delta_{sd}=\Delta_{c}$ for comparison.
@@
where the Hamiltonian takes the form as
${}\hat{\tilde{H}}_{sd}=\Delta_{sd}\hat{a}^{\dagger}\hat{a}+\omega_{m}\hat{b}^{\dagger}\hat{b}+(G_{sd}\hat{a}^{\dagger}+G_{sd}^{*}\hat{a})(\hat{b}^{\dagger}+\hat{b}).$ (16)
@@
Figure 3: The scaled effective coupling strength $G_{sd}$ versus the scaled bare driving detuning $\Delta_{c}/\omega_{m}$ and the scaled Kerr constant $\chi/\omega_{m}$ when the cavity-field decay rate $\kappa_{a}$ takes different values: (a) $\kappa_{a}/\omega_{m}=0.5$, (b) $\kappa_{a}/\omega_{m}=0.8$, (c) $\kappa_{a}/\omega_{m}=1.2$, and (d) $\kappa_{a}/\omega_{m}=1.5$. Other parameters are the same as Fig. 2.
@@
For generating optomechanical entanglement, both the effective detuning $\Delta_{sd}$ and the effective coupling strength $G_{sd}$ are important parameters. Based on Eqs. (5), (7), (11), and (12), we can see that both $\Delta_{sd}$ and $G_{sd}$ depend on the displacement parameters and the squeezing parameters, which are determined by the system parameters. In Fig. 2, we show the dependence of the normalized detuning $\Delta_{sd}/\omega_{m}$ on $\Delta_{c}/\omega_{m}$ and $\chi/\omega_{m}$ for different values of $\kappa_{a}/\omega_{m}$ in both the sideband resolved and unsolved regimes. Note that Fig. 2 is plotted only for parameter sets satisfying the steady-state condition, and that the blank area represents the region where the steady-state condition is not met [see Sec. IV(A) for details]. We can see from Figs. 2(a)-2(d) that the effective driving detuning $\Delta_{sd}/\omega_{m}$ remains positive for these values of $\kappa_{a}/\omega_{m}$, indicating that the normalized detuning $\Delta_{sd}/\omega_{m}$ caused by both the displacement and squeezing transformations will move towards the red sideband direction. Namely, $\Delta_{sd}/\omega_{m}$ becomes positive within some parameter regions of $\Delta_{c}<0$. For a fixed $\Delta_{c}/\omega_{m}$, the normalized detuning difference $(\Delta_{sd}-\Delta_{c})/\omega_{m}$ increases with the increase of the parameter $\chi$ (as shown by the difference between the surfaces of $\Delta_{sd}/\omega_{m}$ and $\Delta_{c}/\omega_{m}$). In addition, we find that the surface patterns of $\Delta_{sd}/\omega_{m}$ corresponding to the four cases of $\kappa_{a}/\omega_{m}$ exhibit slightly difference. We also find that the red-sideband resonance $\Delta_{sd}/\omega_{m}=1$ corresponds to a relatively small $\chi/\omega_{m}$, and the intersection lines associated with $\Delta_{sd}/\omega_{m}=1$ have similar trend for the four cases of $\kappa_{a}/\omega_{m}$.
@@ IV.1 Analyses of the stability of the system
For the linearized optomechanical system, its steady state can be characterized by the Gaussian state. Since the equation of motions for the displacement amplitudes $\alpha(t)$ and $\beta(t)$ are nonlinear, there may exist rich stability in this system. In our following study of the optomechanical entanglement, we mainly focus on the single stable regime. Therefore, we choose the parameters such that the steady-state displacement amplitudes $\alpha_{ss}$ and $\beta_{ss}$ have a single value. Meanwhile, we use the Routh-Hurwitz criterion to confirm that this steady state solution is stable. The solution of the amplitudes $\alpha_{ss}$ and $\beta_{ss}$ is obtained by solving Eqs. (5). To further examine the stability around this working point, we need to evaluate the equation of motion for the first-order moments. To this end, we derive the Heisenberg equation for $\hat{a}$ and $\hat{b}$ based on the linearized optomechanical Hamiltonian $\hat{\tilde{H}}_{sd}$. To include the dissipations and functuations, we phenomenologically add both the dissipation term and the noise operators into the equations of motion, then we obtain the following Langevin equation,
@@
For convenience, we work in the quadrature representation by introducing the quadrature operators, $\hat{X}_{o}=(\hat{o}^{\dagger}+\hat{o})/\sqrt{2}$ and $\hat{Y}_{o}=i(\hat{o}^{\dagger}-\hat{o})/\sqrt{2}$ for $\hat{o}=\hat{a}$ and $\hat{b}$. In the quadrature representation, we denote $\hat{\mathbf{u}}(t)=(\hat{X}_{a},\hat{Y}_{a},\hat{X}_{b},\hat{Y}_{b})^{T}$ (“$T$” denotes the matrix transpose), then the Langevin equations can be expressed as
@@
The stability of the system can be analyzed by checking the real parts of all the eigenvalues of the coefficient matrix $\mathbf{A}$. If all the real parts of the eigenvalues are negative, then the system is stable. It can be seen from Eqs. (5) that the equation could have three solutions under certain parameter conditions, corresponding to three fixed points. If all the three fixed points satisfy the steady-state condition, the system enters a tristable regime. If only two fixed points satisfy the steady-state condition, the system is bistable. This implies that the presence of three solutions may lead the system into a multi-stable regime. Therefore, we exclude the case with three solutions and focus on the single-solution scenario. The first step for clarifying the single stable solution is to distinguish between the multi-solution and single-solution regions. The single-solution cases are then classified as either steady or non-steady using the Routh-Hurwitz criterion.

### v2

# Kerr enhanced optomechanical entanglement generation via reservoir design

@@ Abstract
Quantum entanglement is a crucial resource in quantum technologies, enabling advancements in quantum computing, quantum communication, and quantum precision measurement. Here, we propose a method to enhance optomechanical entanglement by introducing an optical Kerr nonlinear medium and a squeezed vacuum reservoir of the optomechanical cavity. By performing the displacement and squeezing transformations, the system can be reduced to a standard linearized optomechanical system with normalized driving detuning and linearized-coupling strength, in which the optical and mechanical modes are, respectively, coupled to an optical vacuum bath and a mechanical heat bath. We focus on the entanglement generation in the single stable regime of the system. By evaluating the steady-state logarithm negativity, we find that the optomechanical entanglement can be enhanced within a wide range of the Kerr constant. In addition, the Kerr nonlinearity can extend the stable region, enabling considerable entanglement generation in the blue-sideband parameter region. We also investigate the dependence of the entanglement generation on the average thermal phonon occupation of the mechanical bath and the optical driving amplitude. It is found that the presence of the Kerr nonlinearity allows the generation of optomechanical entanglement even when the thermal phonon occupation of the mechanical bath is as high as 3000. Our findings will provide valuable insights into enhancing fragile quantum resources in quantum systems.
@@ I Introduction
Cavity optomechanics primarily investigates the radiation-pressure interaction between optical modes and mechanical modes, along with the novel quantum effects and applications arising from this interaction [1, 2, 3, 4]. The study of cavity optomechanics holds dual significance in both fundamental theoretical research [5, 6] and cutting-edge applied research [7]. Not only does it advance the fundamental theoretical exploration of quantum theory, including macroscopic quantum behaviors [8, 9, 10] and quantum-classical boundary [5, 11], but it also advances the applied research value in frontier quantum technologies, especially concerning quantum precision measurement [12, 13, 14] and quantum sensing [15, 16, 17]. In addition, several experiments in cavity optomechanics have demonstrated ultrasensitive detection capabilities near the classical or standard quantum limits [18, 19, 20]. In particular, quantum entanglement in a cavity optomechanical system holds significant research meaning for all these aspects [21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33], because quantum entanglement is located at the kernel position of quantum theory [34, 35], and it is an important resource in quantum information science [36]. Note that quantum entanglement in other physical systems has also been investigated in recent years [37, 38, 39, 40, 41].
In general, it is expected to generate strong entanglement for implementing various quantum information processing tasks. In cavity optomechanics, various strategies have been proposed, such as reservoir engineering techniques [42, 43, 44], multimode coupling and collective effects [45, 46, 47], the injection of squeezed light [48], and the quantum interference effect [49, 50]. Usually, quantum entanglement induced by interaction can be enhanced via increasing the coupling strength. For linearized optomechanical interaction, the coupling strength can be enhanced by increasing the driving strength. However, the large driving amplitude is subject to the unstability issue [51]. This motivates us to explore other methods of enhancing the optomechanical interaction.
@@ II Physical model
Figure 1: Schematic of the Kerr-cavity optomechanical system consisting of a mechanical mode $b$ optomechanically coupled to a optical mode $a$ containing a Kerr nonlinear medium with nonlinear constant $\chi$. The $g_{0}$ is the single-photon optomechancial coupling strength. The cavity field and the mechanical resonator are, respectively, contacted to a squeezed vacuum reservoir (with decay rate $\kappa_{a}$, squeezing parameter $r_{e}$, and reference phase $\theta_{e}$) and a heat bath (with decay rate $\kappa_{b}$ and average thermal phonon occupation $\bar{n}_{th}$). In addition, the cavity field is driven by a monochromatic field with driving amplitude $\Omega$ and frequency $\omega_{d}$.
The system under consideration is a Kerr-cavity optomechanical system (Fig. 1), which is formed by an optomechanical cavity containing a Kerr nonlinear medium. We consider the adiabatic regime of the cavity field, and hence focus on a single field mode in the cavity. The cavity field is coupled to the mechanical mode through the radiation-pressure interaction. The cavity is driven by a monochromatic field, and it is connected to a squeezed vacuum bath, while the mechanical mode is coupled to a heat bath. In a rotating frame defined by the unitary transformation operator $\mathrm{exp}(-i\omega_{d}\hat{a}^{{\dagger}}\hat{a}t)$, with $\omega_{d}$ being the driving frequency, the Hamiltonian of the Kerr cavity optomechanical system reads ($\hbar=1$)
[…]
@@
To include the dissipation in the system, we assume that the cavity-field mode is connected to a squeezed vacuum reservoir with the central frequency $\omega_{d}$, and the mechanical mode is coupled to a heat bath. Here, the squeezed vacuum reservoir is introduced to counteract the thermal noise caused by the two-photon terms, which is induced by the Kerr nonlinearity under the linearization. In the Markovian-dissipation regime, the evolution of the system is governed by the quantum master equation
@@ III Enhanced optomechanical interaction
[…]
@@
In the derivation of Eq. (III), we have eliminated the linear term of the operators $\hat{a}$ ($\hat{a}^{\dagger}$) and $\hat{b}$ ($\hat{b}^{\dagger}$) by setting their coefficients to be zero. In this way, we obtain the equations of motion of the displacement amplitudes $\alpha(t)$ and $\beta(t)$ as follows:
@@
Here, $\hat{\rho}_{sd}$ is the density matrix of the system in the squeezed representation and $\hat{S}_{a}(\xi)=\exp[(\xi^{\ast}\hat{a}^{2}-\xi\hat{a}^{\dagger 2})/2]$ is the squeezing operator with $\xi=r\exp(i\varphi)$, where $r$ is the squeezing parameter and $\varphi$ is the reference phase for the squeezed field. In the squeezed representation, the quantum master equation is given by
@@
where $\alpha_{r}$ and $\alpha_{i}$ are, respectively, the real and imaginary parts of $\alpha_{ss}$, i.e., $\alpha_{ss}=\alpha_{r}+i\alpha_{i}$.
@@
$\!\times(\sinh r_{e}\sinh r\!+e^{i(\varphi\!-\theta_{e})}\cosh r_{e}\cosh r).$ (14b)
@@
Below, we analyze how to obtain the solution of $r_{e}$ and $\theta_{e}$. In principle, under the given $r$ and $\varphi$, we can analyze the parameters $N_{ss}$ and $M_{ss}$ as functions of $\theta_{e}$ and $r_{e}$. By analyzing Eqs. (14), we find that $N_{ss}=0$ and $M_{ss}=0$ when $\theta_{e}=\varphi+n\pi$ $(n=1,3,5,\dots)$ and $r_{e}=r$. It should be emphasized that $N_{ss}=0$ and $M_{ss}=0$ imply that both the thermal noise and the squeezing vacuum noise have been completely suppressed. This working point is very important to observe the optomechanical effects at the single-photon level [52]. This feature is also the motivation for introducing the squeezing vacuum reservoir, i.e., using a well-designed squeezing vacuum bath to suppress the thermal noise and excitations caused by the two-photon driving.
Figure 2: The normalized detuning $\Delta_{sd}/\omega_{m}$ vs the scaled bare driving detuning $\Delta_{c}/\omega_{m}$ and the scaled Kerr constant $\chi/\omega_{m}$ when the cavity-field decay rate $\kappa_{a}$ takes different values: (a) $\kappa_{a}/\omega_{m}=0.5$, (b) $\kappa_{a}/\omega_{m}=0.8$, (c) $\kappa_{a}/\omega_{m}=1.2$, and (d) $\kappa_{a}/\omega_{m}=1.5$. Other parameters used are $g_{0}/\omega_{m}=0.005$, $\kappa_{b}/\omega_{m}=10^{-5}$, $\Omega/\omega_{m}=50$, and $\bar{n}_{th}=0$. Here, we also show the surfaces corresponding to $\Delta_{sd}/\omega_{m}=1$ and $\Delta_{sd}=\Delta_{c}$ for comparison.
@@
where the Hamiltonian takes the form
[…]
@@
Although the Hamiltonian in Eq. (16) takes the form of a standard linearized optomechanical Hamiltonian, it is important to emphasize that the effective parameters $\Delta_{sd}$ and $G_{sd}$ are nontrivially modified by the Kerr nonlinearity and the squeezed vacuum reservoir. The displacement and squeezing transformations reshape the structure of the parameter space, enabling the system to access a new working point of the larger optomechanical entanglement. Therefore, our method does not simply reproduce a standard linearized optomechanical system but rather provides an experimentally feasible route to expand the entanglement-enhancing parameter regime while ensuring dynamical stability.
We should point out that, in the present scheme, the entangling power [68] associated with the linearized optomechanical interaction will not be enhanced by introducing the Kerr nonlinearity and the squeezed vacuum bath. This is because the quantum master equation (15) describing the transformed optomechanical system takes the same form as that of a standard linearized optomechanical system. Nevertheless, at certain working points, the entanglement generation can be enhanced because an optimal working point can be achieved owing to the nonlinear mapping of the parameter space. In addition, a further study of the entangling power for the cavity optomechanical model deserves to be conducted because it can give the real upper bound of optomechanical entanglement creation, which will be very important for exploiting quantum technology based on the cavity optomechanical platform.
@@
Figure 3: The scaled effective coupling strength $G_{sd}$ vs the scaled bare driving detuning $\Delta_{c}/\omega_{m}$ and the scaled Kerr constant $\chi/\omega_{m}$ when the cavity-field decay rate $\kappa_{a}$ takes different values: (a) $\kappa_{a}/\omega_{m}=0.5$, (b) $\kappa_{a}/\omega_{m}=0.8$, (c) $\kappa_{a}/\omega_{m}=1.2$, and (d) $\kappa_{a}/\omega_{m}=1.5$. Other parameters are the same as Fig. 2.
@@
For generating optomechanical entanglement, both the effective detuning $\Delta_{sd}$ and the effective coupling strength $G_{sd}$ are important parameters. Based on Eqs. (5), (7), (11), and (12), we can see that both $\Delta_{sd}$ and $G_{sd}$ depend on the displacement parameters and the squeezing parameters, which are determined by the system parameters. In Fig. 2, we show the dependence of the normalized detuning $\Delta_{sd}/\omega_{m}$ on $\Delta_{c}/\omega_{m}$ and $\chi/\omega_{m}$ for different values of $\kappa_{a}/\omega_{m}$ in both the sideband resolved and unsolved regimes. Note that Fig. 2 is plotted only for parameter sets satisfying the steady-state condition, and that the blank area represents the region where the steady-state condition is not met (see Sec. IVA for details). We can see from Figs. 2(a)-2(d) that the effective driving detuning $\Delta_{sd}/\omega_{m}$ remains positive for these values of $\kappa_{a}/\omega_{m}$, indicating that the normalized detuning $\Delta_{sd}/\omega_{m}$ caused by both the displacement and squeezing transformations will move towards the red sideband direction. Namely, $\Delta_{sd}/\omega_{m}$ becomes positive within some parameter regions of $\Delta_{c}<0$. For a fixed $\Delta_{c}/\omega_{m}$, the normalized detuning difference $(\Delta_{sd}-\Delta_{c})/\omega_{m}$ increases with the increase of the parameter $\chi$ (as shown by the difference between the surfaces of $\Delta_{sd}/\omega_{m}$ and $\Delta_{c}/\omega_{m}$). In addition, we find that the surface patterns of $\Delta_{sd}/\omega_{m}$ corresponding to the four cases of $\kappa_{a}/\omega_{m}$ exhibit slight difference. We also find that the red-sideband resonance $\Delta_{sd}/\omega_{m}=1$ corresponds to a relatively small $\chi/\omega_{m}$, and the intersection lines associated with $\Delta_{sd}/\omega_{m}=1$ have similar trend for the four cases of $\kappa_{a}/\omega_{m}$.
@@ IV.1 Analyses of the stability of the system
For the linearized optomechanical system, its steady state can be characterized by the Gaussian state. Since the equation of motion for the displacement amplitudes $\alpha(t)$ and $\beta(t)$ are nonlinear, rich stability may exist in this system. In our following study of the optomechanical entanglement, we mainly focus on the single stable regime. Therefore, we choose the parameters such that the steady-state displacement amplitudes $\alpha_{ss}$ and $\beta_{ss}$ have a single value. Meanwhile, we use the Routh-Hurwitz criterion to confirm that this steady-state solution is stable. The solution of the amplitudes $\alpha_{ss}$ and $\beta_{ss}$ is obtained by solving Eqs. (5). To further examine the stability around this working point, we need to evaluate the equation of motion for the first-order moments. To this end, we derive the Heisenberg equation for $\hat{a}$ and $\hat{b}$ based on the linearized optomechanical Hamiltonian $\hat{\tilde{H}}_{sd}$. To include the dissipations and fluctuations, we phenomenologically add both the dissipation term and the noise operators into the equations of motion, then we obtain the following Langevin equation,
@@
For convenience, we work in the quadrature representation by introducing the quadrature operators, $\hat{X}_{o}=(\hat{o}^{\dagger}+\hat{o})/\sqrt{2}$ and $\hat{Y}_{o}=i(\hat{o}^{\dagger}-\hat{o})/\sqrt{2}$ for $\hat{o}=\hat{a}$ and $\hat{b}$. In the quadrature representation, we denote $\hat{\mathbf{u}}(t)=(\hat{X}_{a},\hat{Y}_{a},\hat{X}_{b},\hat{Y}_{b})^{T}$ (“TT” denotes the matrix transpose); then the Langevin equations can be expressed as
@@
The stability of the system can be analyzed by checking the real parts of all the eigenvalues of the coefficient matrix $\mathbf{A}$. If all the real parts of the eigenvalues are negative, then the system is stable. It can be seen from Eqs. (5) that the equation could have three solutions under certain parameter conditions, corresponding to three fixed points. If all three fixed points satisfy the steady-state condition, the system enters a tristable regime. If only two fixed points satisfy the steady-state condition, the system is bistable. This implies that the presence of three solutions may lead the system into a multistable regime. Therefore, we exclude the case with three solutions and focus on the single-solution scenario. The first step for clarifying the single stable solution is to distinguish between the multi-solution and single-solution regions. The single-solution cases are then classified as either steady or nonsteady using the Routh-Hurwitz criterion.